# Succeeding at Scale: Automated Multi-Retriever Fusion and Query-Side Adaptation for Multi-Tenant Search

### v1

@@ 5 Conclusion
We presented a unified approach to scalable dataset construction and efficient model adaptation for multi-tenant retrieval. Our automated pipeline, combining multi-retriever fusion with LLM-based consistency filtering, produces high-quality training data without manual annotation, yielding the DevRev Search benchmark. Our index-preserving adaptation strategy, fine-tuning only the query encoder via LoRA, achieves competitive performance while eliminating re-indexing overhead, with layer-targeted updates further optimizing the quality-efficiency trade-off.
Limitations
Our evaluation is limited to English queries across two domains; generalizability to other verticals and multilingual settings remains unexplored. Query-side adaptation may impose a performance ceiling compared to joint fine-tuning, and we do not investigate hybrid approaches with cross-encoder rerankers. Optimal LoRA configurations may vary across architectures, and our metrics do not directly measure downstream RAG performance.
Acknowledgments
We used AI writing assistants to assist with code development and manuscript refinement. All content was reviewed and verified by the authors.
Appendix
Appendix A Dataset Details
A.1 Document Segmentation and Semantic Granularity
Enterprise documentation is often characterized by high length and low information density relative to specific user intents. Large documents pose two primary challenges for dense retrieval: (1) Context Constraints, as most transformer-based encoders have a fixed token limit (e.g., 512 tokens), and (2) Information Compression, where the fixed-dimensional latent vector is insufficient to represent the entire semantic breadth of a long document, leading to "diluted" embeddings.
To mitigate these issues, we collected all public-facing documentation from DevRev and applied a Recursive Character Splitting strategy LangChain (2025). We chose a maximum chunk size of 500 characters with zero overlap to maximize the number of distinct semantic units. The recursive algorithm prioritizes splitting at logical boundaries (e.g., double newlines, then periods, then spaces) to maintain structural integrity. This ensures that each document fragment is small enough to be represented accurately by a single embedding while remaining large enough to contain a self-contained answer or concept.
A.2 Dataset Statistics
@@
We partition the DevRev Search dataset into a training set and an evaluation (test) set. The dataset consists of 291 training queries and 92 test queries. The training set exhibits a rich density of relevant documents, with an average of $13.61$ golden chunks per query. Notably, the distribution of relevant documents is right-skewed, with a median of 6 and a high standard deviation ($\sigma=21.41$). This variance reflects the diverse nature of enterprise search, where some queries address specific technical identifiers with a single relevant source, while others address broad architectural topics with many relevant documentation fragments. To maintain the integrity of our public benchmark, we withhold the gold labels for the test queries to facilitate blind evaluation.
A.3 Design Rationale
Our pipeline reflects a principled balance between recall-oriented aggregation and precision-oriented filtering. The multi-retriever ensemble, comprising seven distinct models, ensures comprehensive coverage across the document corpus. By combining dense semantic encoders with lexical models (BM25), we capture complementary aspects of relevance that a single model might overlook. Following this "wide net" approach, LLM-based filtering leverages deep reasoning capabilities to adjudicate the final labels, eliminating false positives that exhibit surface-level similarity but fail to satisfy the semantic requirements of the query.
A.4 Analysis of Retriever Contributions
To validate the necessity of a multi-retriever ensemble, we investigate the individual contributions and potential redundancies of each model.
Individual Recall.
We first evaluate the standalone performance of each retriever. For a given query, we retrieve the top 420 candidates from a single model and calculate its recall against the final generated ground truth. As shown in Table 1, even the highest-performing model (gemini-embedding-001) achieves only 82.48% recall. This confirms that relying on a single retriever—no matter how powerful—would result in a significant loss of valid relevant documents, thereby biasing the dataset.
Redundancy and Diversity.
We further conduct a "leave-one-out" ablation study to determine if any retriever is redundant. For each model $M_{i}$, we aggregate 60 candidates from each of the other six retrievers (as we did during dataset generation) and measure the union’s recall against the ground truth. If the recall for a combination $\text{All}\setminus\{M_{i}\}$ were to reach 1.0, it would indicate that $M_{i}$ is redundant.As reported in Table 2, no model combination achieves perfect recall, with values ranging from 93.25% to 97.13%. This signifies that every retriever in our ensemble contributes unique relevant candidates that are not captured by the others. Notably, the drop in recall is most significant when removing the top-performing dense models, but even the removal of BM25 results in a loss of coverage, underscoring the necessity of a hybrid, multi-model approach for high-quality dataset construction.
@@ A.4 Analysis of Retriever Contributions
Table 1: Recall@420 of Individual Retrievers on DevRev Search dataset
@@
Table 2: Leave-one-out ablation study on DevRev Search dataset
A.5 LLM-based Filtering Prompt
@@ A.5.2 Few-Shot Examples
A.5.3 User Prompt Template
@@
A.6 Experimental Setup
We use AdamW optimizer with cosine learning rate scheduler.
Table 3 shows the hyperparameter configuration used by us.
Dataset Model Learning Rate Warmup Steps (%) LoRA Dropout (if LoRA)
DevRev Search snowflake-arctic-mbed-l-v2 $5\mathrm{e}{-6}$ 0 0.00
DevRev Search Qwen3-Embedding-4B $5\mathrm{e}{-6}$ 1010 0.05
Scifact snowflake-arctic-mbed-l-v2 $5\mathrm{e}{-6}$ 0 0.00
Scifact Qwen3-Embedding-4B $5\mathrm{e}{-6}$ 1010 0.05
Table 3: Hyperparameter settings for different datasets and embedding models

### v2

@@ 5 Conclusion
We presented a unified approach to scalable dataset construction and efficient model adaptation for multi-tenant retrieval. Our automated pipeline, combining multi-retriever fusion with LLM-based consistency filtering, produces high-quality training data without manual annotation, yielding the DevRev-Search benchmark. Our index-preserving adaptation strategy, fine-tuning only the query encoder via PEFT, achieves competitive performance while eliminating re-indexing overhead, thus optimizing for the quality-efficiency trade-off.
Appendix A Appendix
A.1 Dataset Statistics
@@
We partition the DevRev-Search dataset into a training set and an evaluation (test) set. The dataset consists of 291 training queries and 92 test queries. The training set exhibits a rich density of relevant documents, with an average of $13.61$ golden chunks per query. Notably, the distribution of relevant documents is right-skewed, with a median of 6 and a high standard deviation ($\sigma=21.41$). This variance reflects the diverse nature of enterprise search, where some queries address specific technical identifiers with a single relevant source, while others address broad architectural topics with many relevant documentation fragments. To maintain the integrity of our public benchmark, we withhold the gold labels for the test queries to facilitate blind evaluation.
A.2 Retriever Contributions
Table 3: Recall@420 of Individual Retrievers on DevRev-Search dataset
@@ A.2 Retriever Contributions
Table 4: Leave-one-out ablation study on DevRev-Search dataset
@@
A.3 LLM-based Filtering Prompt
@@ A.3.2 Few-Shot Examples
A.3.3 User Prompt Template
@@
A.4 Experimental Setup
All our trainings optimize InfoNCE loss using AdamW optimizer, with 16 mined hard negatives per query, with 8 negatives randomly sampled from them each time. A learning rate of $5\times 10^{-6}$ is used for all runs with a cosine scheduler and no warm-up. For all the experiments, we use a uniform temperature of 0.1, which was found to work the best.
Additionally, we noticed that simply using fixed mined hard negatives leads to cases of representation collapse wherein InfoNCE loss keeps improving, but training metrics rapidly deteriorate after a point, especially in datasets like FiQA and DevRev-Search. Upon further analysis, we noticed that this happens due to evolving hard negatives, where over the course of training, old hard negatives become ’simple’ and new hard negatives replace them. To counteract this, we adopt the asynchronous ANCE training proposed by Xiong et al. (2020) to periodically update the hard negatives in InfoNCE loss at every 200 steps. This proved to improve the performance in most cases, in addition to much more stable training as shown in Fig 2. Quantitative comparison of results obtained with and without ANCE is shown in Table 6
The feed-forward network (FFN) used in our PEFT experiments consists of GELU (Hendrycks (2016)) activations and 2 hidden layers, each with the same number of nodes as the base model’s embedding dimension. Also, for both linear and FFN heads, we find that zeroing biases and initializing weight matrices to the identity perform consistently better than standard initializations such as Xavier (Glorot and Bengio (2010)) or Kaiming (He et al. (2015)).
Table 5: Parameter efficiency comparison for Query-Only adaptation across LoRA ranks, unfrozen layers, and head-only training.
Model Configuration Trainable Total % Full
Params Params
arctic-l-v2 Full Model 566,705,152 566,705,152 100.00
LoRA $r=8$ 3,538,944 570,244,096 0.62
LoRA $r=16$ 7,077,888 573,783,040 1.25
LoRA $r=32$ 14,155,776 580,860,928 2.50
LoRA $r=64$ 28,311,552 595,016,704 4.99
LoRA $r=128$ 56,623,104 623,328,256 9.99
Linear Head (Frozen Base) 1,049,600 567,754,752 0.19
MLP Head (Frozen Base) 3,148,800 569,853,952 0.56
1 Unfrozen Layer 12,596,224 566,705,152 2.22
2 Unfrozen Layers 25,192,448 566,705,152 4.45
4 Unfrozen Layers 50,384,896 566,705,152 8.89
8 Unfrozen Layers 100,769,792 566,705,152 17.78
qwen3-4b Full Model 4,021,774,336 4,021,774,336 100.00
LoRA $r=8$ 16,515,072 4,038,289,408 0.41
LoRA $r=16$ 33,030,144 4,054,804,480 0.82
LoRA $r=32$ 66,060,288 4,087,834,624 1.64
LoRA $r=64$ 132,120,576 4,153,894,912 3.29
LoRA $r=128$ 264,241,152 4,286,015,488 6.57
Linear Head (Frozen Base) 6,556,160 4,028,330,496 0.16
MLP Head (Frozen Base) 19,668,480 4,041,442,816 0.49
1 Unfrozen Layer 100,933,376 4,021,774,336 2.51
2 Unfrozen Layers 201,864,192 4,021,774,336 5.02
4 Unfrozen Layers 403,725,824 4,021,774,336 10.04
8 Unfrozen Layers 807,449,088 4,021,774,336 20.08
Table 6: Retrieval performance with and without ANCE for full Query-Only adaptation
Model DevRev-Search SciFact FiQA
recall@10 ndcg@10 recall@10 ndcg@10 recall@10 ndcg@10
arctic-l-v2 Standard 0.291 0.331 0.860 0.731 0.543 0.473
+ANCE 0.296 0.343 0.854 0.727 0.554 0.478
qwen3-4b Standard 0.333 0.373 0.941 0.813 0.679 0.593
+ANCE 0.327 0.370 0.953 0.830 0.688 0.602
Figure 2: Training NDCG@10 vs training step for Query-Only adaptation of arctic-l-v2 (full model)
A.5 Ablation Results
Table 7: Retrieval performance variation across different LoRA ranks ($r=\alpha$ for all configurations)
Model LoRA Rank DevRev-Search SciFact FiQA
recall@10 ndcg@10 recall@10 ndcg@10 recall@10 ndcg@10
arctic-l-v2 8 0.258 0.307 0.867 0.738 0.540 0.469
16 0.268 0.322 0.880 0.748 0.548 0.474
32 0.283 0.328 0.905 0.769 0.554 0.478
64 0.304 0.341 0.914 0.807 0.555 0.478
128 0.309 0.342 0.915 0.827 0.555 0.478
qwen3-4b 8 0.304 0.346 0.948 0.826 0.685 0.603
16 0.329 0.360 0.948 0.832 0.686 0.605
32 0.355 0.399 0.950 0.844 0.690 0.607
64 0.336 0.369 0.950 0.816 0.694 0.607
128 0.336 0.380 0.957 0.831 0.682 0.601
Table 8: Retrieval performance variation across different target modules for LoRA $r=64$ and $\alpha=64$
Model Layers DevRev-Search SciFact FiQA
recall@10 ndcg@10 recall@10 ndcg@10 recall@10 ndcg@10
arctic-l-v2 Dense 0.287 0.329 0.854 0.726 0.551 0.477
QV 0.283 0.325 0.854 0.726 0.554 0.476
QKV 0.283 0.325 0.854 0.726 0.554 0.477
All 0.304 0.341 0.914 0.807 0.555 0.478
qwen3-4b Dense 0.333 0.371 0.950 0.818 0.693 0.607
QV 0.324 0.377 0.923 0.793 0.688 0.605
QKV 0.337 0.368 0.930 0.796 0.689 0.606
All 0.336 0.369 0.950 0.816 0.694 0.607
Table 9: Retrieval performance variation across different numbers of unfrozen transformer layers at the top of the base model, with the rest of the layers frozen
Model Layers DevRev-Search SciFact FiQA
recall@10 ndcg@10 recall@10 ndcg@10 recall@10 ndcg@10
arctic-l-v2 1 0.258 0.306 0.827 0.700 0.534 0.462
2 0.260 0.306 0.835 0.709 0.536 0.464
4 0.267 0.311 0.835 0.712 0.539 0.471
8 0.273 0.317 0.849 0.718 0.547 0.476
qwen3-4b 1 0.237 0.290 0.903 0.778 0.661 0.576
2 0.259 0.306 0.908 0.783 0.664 0.579
4 0.275 0.325 0.912 0.788 0.667 0.583
8 0.312 0.356 0.932 0.802 0.673 0.590